Joes team can completely manicure the lawn at the Detroit Tiger's Comerica Park in 12 hours. When Sally's team was brought on to help, the job was done in about 5 1/2 hours. How long does it take Sally's team to completely manicure the lawn when they are working alone?
Answer by checkley77(12569)

12X/(12+X)=5.5 CROSS MULTIPLY
12X=5.5(12+X)
12X=66+5.5X
12X-5.5X=66
6.5X=66
X=66/6.5
X=10.1538 OR 10 HRS. 9 MIN. FOR SALLY'S TEAM WHEN WORKING ALONE.
12*10.1538/(12+10.1538)=5.5
121.8456/22.1538=5.5
5.5~5.5